# Evidence of Scaling Regimes in the Hopfield Dynamics of Whole Brain Model

### v1

@@
Abstract
@@
It is shown that a Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently introduced by Deco Deco2020 , according to which the process of information transfer within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows. Although both models employ a coupling strength which decays exponentially with the euclidean distance between the nodes, their mathematical nature is widely different, Hopf oscillators versus Hopfield neural network. Hence, their convergence suggests a remarkable robustness of the aforementioned scaling picture. Furthermore, the present analysis shows that the Hopfield model brain remains functional by removing links above about five decay lengths, corresponding to about one sixth of the size of the global brain. This suggests that, in terms of connectivity decay length, the Hopfield brain functions in a sort of intermediate “turbulent liquid”-like state, whose essential connections are the intermediate ones between the connectivity decay length and the global brain size. This ”turbulent-like liquid” appears to be more spiky than actual turbulent fluids, with a scaling exponent around $2/5$ instead of $2/3$.
I Introduction
@@
In a recent paper Deco2020 , it was argued that the information transfer within the human brain may proceed in close analogy with the mechanisms which govern mass and energy transport in turbulent fluids. The analogy refers to the emergence of correlated spacetime patterns which would facilitate and possibly optimize information transfer across the brain, the “same” way as fluid turbulence enhances energy transport across fluids. Based on an extended dynamical system of oscillators (Hopf model), supplemented by large-scale neuro-imaging empirical data on the connections between different brain areas, the authors of Deco2020 numerically showed that the statistical correlations of human brain signals may exhibit a scaling regime similar to the one observed in turbulent fluids.
@@
At a closer scrutiny, though, the analogy does not appear to be quantitative, meaning that the spatial scaling exponents of the second order structure function $S_{2}(d)=<(O(d)-O(0))^{2}>\sim d^{\alpha}$, $O$ being a suitable order parameter, are not the same, about $\alpha\sim 1/2$ for the Hopf model of the human brain against $\alpha$=2/3 for fluid turbulence. Yet, a scaling regime in space is found in either cases, which is intriguing per se, regardless of the specific value of $\alpha$.
@@
In this paper, we probe the robustness of the above analogy by performing a similar analysis based on the same neuro-data, but using a very different model, namely a Hopfield neural network, and measuring scaling exponents as a function of the connectivity decay length $\delta$.
@@
The recurrent Hopfield neural network model is a well-established model for small-scale neuronal networks Amari1972 ; Little1974 ; Hopfield1982 . It is a neural network model based on binary McCulloch-Pitts neurons Amit1985 ; Brunel2016 ; Hillar2018 ; Hillar2021 in which every processing unit ($i$) is connected to all the other ones ($j$) through a set of weights ($J_{ij}$). Most of the literature considers fully connected and symmetric Hopfield networks since this simplifies the mathematical analysis of the model, although recently the investigation of diluted networks Stauffer2003 ; Brunel2016 ; Kim2017 ; Folli2018 ; Leonetti2020 and asymmetric networks Gopalsamy1994 ; Xu1996 ; Chen2001 ; Franca2000 ; Zheng2010 ; Szedlak2014 ; Folli2018 ; Leonetti2020 was introduced.
@@
Usually, a recurrent Hopfield neural network is trained to store patterns or memories as steady state attractors with the Hebbian prescription. In this case, the coupling matrix $J_{ij}$ assumes the form of the sum of diadic forms, it is therefore symmetric and fully connected. The maximum storage capacity of such a network is found to be linear with the number of nodes $N$. The capacity can be increased by the process of learning (Hebbian learning): in this process the simultaneous activation of two nodes will strengthen their connection. This is inspired by biological neural networks where the simultaneous activation of neurons leads to increments in synaptic strength Hebb1949 , and it allows the network to store a larger number of patterns Amit1985 , but this quantity still remains linear with $N$. Lately, researchers have also investigated how unlearning methods can improve the stability of the stored patterns Fachechi2019 ; Benedetti2022 ; Benedetti2023 .
@@
Interestingly, it has been recently demonstrated both numerically Folli2018 ; Gosti2019 and theoretically Hwang2019 ; Hwang2020 that random, non-Hebbian, coupling matrices have a storage capacity much larger than the Hebbian ones. Indeed the storage capacity $C$ of random matrices scales exponentially with $N$: $C\sim e^{\Sigma N}$ being $\Sigma$ the complexity. The latter quantity is found to be maximised for asymmetric and diluted matrices, with values that match the connectivity found in the hippocampus areas.
@@
The approach outlined in reference Deco2020 , as elaborated in the subsequent chapter, diverges from both the Hebbian and random connectivity models. It relies on a deterministic coupling matrix, which is derived from experimental observations of the connections between human brain regions. This approach assumes that the coupling strength is proportionally linked to a specific function of the distances between nodes (see Eq. 3): $J_{ij}=e^{-d_{ij}/\delta}$, $d_{ij}$ being the Euclidean distance between nodes $i$ and $j$ in the network. At variance with the latter approach, where the dynamics is based on the Hopf oscillator model, we here use the same coupling matrix of Deco2020 , but employ the Hopfield dynamics that, to the best of our knowledge, was never used to model whole brain activity dynamics.
@@
ii) At $\delta\sim 5.5$ mm we found $\alpha\sim 2/5$, very close to the one reported by Deco et al. Deco2020 for the same value of $\delta$  [^1] In Ref. Deco2020 the used parameter is the inverse of the decay scale, $\lambda=1/\delta$..
iii) The scaling exponent shows a decreasing trend with the decay length $\delta$. Importantly, the functional relation $\alpha(\delta)$ shows a steep increase in the region around $\delta\sim 5.55$ mm, the value indicated in Deco2020 as the physiologically relevant one. This means that the specific value of the decay length has a significant impact on the scaling exponent, for instance a slight increase to $\delta=6$ mm would yield a scaling exponent slightly above $2/3$, the value associated to homogenoeus incompressible turbulence. Conversely, a slight decrease to $\delta=5$ mm, would yield $\alpha\sim 0.1$, a much less organized regime. This pinpoints the importance of fine-tuning the value of $\delta$.
@@
II The turbulent-like whole-brain model
Deco et al. model the global human brain as a set of $N$ coupled Hopf oscillators of the form
$\frac{dx_{i}}{dt}=a_{i}x_{i}$ $+$ $(\beta y_{i}-x_{i})(x_{i}^{2}+y_{i}^{2})-\omega_{i}y_{i}$ (1)
$+$ $G\sum_{j=1}^{N}J_{ij}(x_{j}-x_{i})-\nu_{i}\xi_{i}$
$\frac{dy_{i}}{dt}=a_{i}y_{i}$ $+$ $(\beta x_{i}+y_{i})(x_{i}^{2}+y_{i}^{2})-\omega_{i}x_{i}$ (2)
$+$ $G\sum_{j=1}^{N}J_{ij}(y_{j}-y_{i})-\nu_{i}\xi_{i}$
In the above, $(x_{i},y_{i})$ are the real and imaginary part of the BOLD signal recorded at the $i$-th node of the network, $a_{i}=-0.02$ are decay terms at the brink of the Hopf bifurcation, where scaling is best observed, $\omega_{n}$ is the natural frequency of the oscillators, $\beta$ the so called shear factor and $\xi$ a gaussian noise with unit variance. The global coupling across the network is represented through the connectivity matrix $J_{ij}$ which is assumed to obey the Exponential Distance Rule
$J_{ij}=e^{-d_{ij}/\delta}$ (3)
where $d_{ij}$ is the euclidean distance between nodes $i$ and $j$ and we have set $\delta$ ($\equiv 1/\lambda$ in Deco’s notation). Deco2020 argues that the exponential distance rule matches very well the empirical HCP dMRI tractography of the human brain. In particular, it reproduces the decay of the fiber densities between the parcel centroid pairs as a function of the Euclidian distance. The authors Deco2020 further show that the optimal fit is obtained for $\delta=5.55$ $\mathrm{m}\mathrm{m}$.
II.1 Amplitude of turbulence and scaling regime
In order to assess the scaling properties of the “turbulent brain” (TB), Deco et al. introduce the following complex Kuramoto Order Parameter:
$\Psi_{i}(t)\doteq R_{i}(t)e^{i\theta_{i}(t)}\doteq\sum_{j=1}^{N}W_{ij}e^{i\phi_{i}(t)}$ (4)
where $W_{ij}=J_{ij}/\sum_{k=1}^{N}J_{ik}$ and $\phi_{i}(t)$ is the phase of the $i$-th oscillator obtain by Hilbert-transform of the (narrow band) BOLD signal. The authors then measure the amplitude of turbulence as the variance of the fluctuations of amplitude of the order parameter, namely:
$D\doteq<R^{2}>-<R>^{2}$ (5)
where bracket indicate average in space and time over the entire network and over hundreds of noise realizations.
The scaling regime is assessed by inspecting the associated structure function
$S_{2}(d)=<(R(x+d)-R(x))^{2}>\sim d^{\alpha}$ (6)
where “$x$” and “$x+d$” indicate that the two network nodes are at distance $d$ and the average is taken at statistical steady state, over all network nodes and noise realizations. This observable is commonly used to assess the statistical properties of fluid turbulence, and for the case of Kolmogorov homogeneous incompressible turbulence one obtains
$\alpha_{T}=2/3,$
indicating that the change in the (longitudinal) flow field at a distance $d$, $\delta u(d)=\langle|u(x+d)-u(x)|\rangle$, scales like $d^{1/3},$ hence the turbulent flow field is not differentiable (sometimes $\alpha$ is also called the roughness exponent) Frisch1995 ; Benzi1993 . For the whole-brain model, Deco et al. Deco2020 report
$\alpha_{D}\sim 1/2,$
indicating that the correlation pattern of the brain signals is less regular than those observed in homogeneous incompressible turbulence.
As we shall show, these values should be taken with extreme caution, as they are very sensitive to a number of assumptions, primarily the structure and functional dependence of the connectivity matrix on the internode distance.
III Discrete-time Hopfield Recurrent Neural Networks
Figure 1: A) 1000 parcel centroids in three dimensional space from Schaefer’s cerebral cortical parcellation atlas Schaefer2018 . B) Frequency histogram of the Euclidean distances $d_{ij}$ between each pair of nodes from Schaefer’s cerebral cortical parcellation atlas Schaefer2018 . The linear dependence on $d_{ij}$ indicates that the nodes are basically distributed on a two dimensional surface. C) Matrix composed of the Euclidean distances $d_{ij}$ between each pair of nodes from Schaefer’s cerebral cortical parcellation atlas Schaefer2018 . D) Matrix formed by the values such that $J_{ij}>0.4$, where $J_{ij}=e^{-d_{ij}/\delta}$ with $\delta$=0.55.
Following Amari1972 ; Little1974 ; Hopfield1982 ; Amit1985 ; Folli2017 ; Gosti2019 ; Leonetti2020 ; Gosti2024 , we consider a network of $N$-binary neurons that can be either firing or silent.
The state of each neuron is a binary variable $s_{i}$, with $i=1,2,\ldots,N$, such that $s_{i}=-1$ or $s_{i}=1$, so that the network state is represented by a binary string $\mathbf{s}=(s_{0},s_{1},\ldots,s_{N})$. Neurons interact through the connectivity matrix $\mathbf{J}$, with matrix elements $J_{ij}$ following the same exponential dependence $J_{ij}=e^{-d_{ij}/\delta}$ used in Deco2020 . We assume a discrete time $t$, and, at each time step $t$, the evolution of the neuron state $s_{i}(t)$ is given by the following non-linear dynamic equation,
@@ III Discrete-time Hopfield Recurrent Neural Networks
where the activation function $sign(x)$ is defined such as $sign(x)=1$ for $x\geq 0$, and $sign(x)=-1$ otherwise. Consequently, whenever the summation of the inputs on neuron $i$ is above zero, the neuron is active (“fires”), otherwise, it remains silent.
III.1 The structural connectivity
We determined the Hopfield Brain connectivity using Schaefer’s cerebral cortical parcellation atlas. This parcellation was obtained from the analysis of a large dataset with a sample size of 1489 Schaefer2018 . This dataset is publicly available, we used Schaefer’s 1000 parcellation with parcel centroids mapped onto the MNI152 volumetric space [^2] Github repository dataset url Schaefer2018 : https://github.com/ThomasYeoLab/CBIG/tree/master/stable_projects/brain_parcellation/Schaefer2018_LocalGlobal/Parcellations/MNI/Centroid_coordinates. Fig. 1A shows the spacial distribution of the centroids in the MNI152 volumetric space. From the centroids of these areas, we computed the pairwise distances. Fig. 1B shows the histogram of the pair distances $d_{ij}$ between each couple of centroids. Figures 1C and D show respectively the matrix composed of the Euclidean distances $d_{ij}$, the matrix obtained with the threshold $J_{ij}>0.4$ were $J_{ij}=e^{-d_{ij}/\delta}$ with $\delta$=0.55.
IV The Structure Factor $S_{2}(d)$
Figure 2: Structure factor $S_{2}(d)$ for $N_{r}=40$ realizations respectively at $\delta=5.55$ mm (A), and $\delta=5.88$ mm (B). The values $S_{2}(d)$ were binned over equal sized intervals. The dashed lines indicate different $\alpha$ values: $\alpha=2/3$ which corresponds to turbulence, $\alpha=1/2$, the value obtained in Deco2020 , and $\alpha=2/5$. We estimate the scaling exponent $\alpha$ with a linear regression of $\log(S(d))\sim\alpha\log(d)+\beta$ in the interval range $[2.7-33.1]$ mm, as used in Deco2020 , and we obtain $\alpha\approx 2/5$ for $\delta=5.55$ mm, and $\alpha\approx 2/3$ for $\delta=5.88$ mm.
We used Deco2020 adaptation of Kolmogorov’s concept of structure functions. In turbulence, the Kolmogorov’s structure factor applies to the longitudinal velocity but in our case it represents the activity of the parcel which as measured through the BOLD signal in a fMRI experiment.
To compute the structure factor we use the auxiliary function $B(d)$ defined as:
$B(d)=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{{\cal{N}}(d)}\sum_{d_{ij}\in d}s_{i}(t)s_{j}(t)$ (8)
@@ IV The Structure Factor $S_{2}(d)$
The function $B(d)$ relates to the structure factor as follows:
$S_{2}(d)$ $\doteq$ $\frac{1}{T}\sum_{t=1}^{T}\frac{1}{{\cal{N}}(d)}\sum_{ij}(s_{i}(t)-s_{j}(t))^{2}\Delta(ij|d)$ (11)
$=$ $2\left[B(0)-B(d)\right]$
@@
The Hopfield dyamics is deterministic, and since the phase space of the nodes’ states is finite, the asymptotic trajectories must necessarily be periodic. Therefore, as also discussed in Folli2017 , Hopfield Recurrent Neural Networks, given an initial state, consistently ends up into a specific steady state (“attractor”). The steady states are deterministic attractors composed of either limit cycles formed by repeating sequence of states, or a single stationary state. Under the parametrizations considered here, we numerically found that the steady states were always composed of a single fixed point, as expected in the case of symmetric couplings $J_{ij}$. In the absence of a further external input (“external stimuli”), the network (the brain) is stuck in the attractor and the whole activity is constant in time, hence no time summation was taken in the previous equations. This is of course a very crude representation of the brain dynamics, which receives continuously “internal” and “external” inputs, but we can think at the whole set of attractors as a reasonable representation of the dynamical correlations within the whole brain model.
@@
Given different values of the decay length $\delta$, we have run the Hopfied model for a system composed by $N=1000$ neurons (with pairwise distances consistent with the Schaefer’s cerebral cortical parcellation atlas Schaefer2018 ). The structure functions have then been calculated by following the system until the fixed points were reached, and by averaging over $N_{r}=40$ randomly chosen initial states. Figure 2 shows the structure factor $S_{2}(d)$ for respectively $\delta=5.55$ mm (Fig. 2A) and $\delta=5.88$ mm (Fig. 2B). The black points with the error bars represent the avarege $S(d)$ computed over 40 different realizations. The dashed green line corresponds to $\alpha=2/3$, the orange line corresponds to $\alpha=1/2$, and the blue line corresponds to $\alpha=2/5$. The $\alpha=1/2$ value (green line) is the value estimated in Deco2020 . The $\alpha=2/3$ is the exponent associated with a turbulece. Due to the significant level of statistical fluctuations, we have estimated the average exponent in two different ways, first as the correct ensemble average $<S_{2}(d)>\sim d^{\alpha}$ and also as an arithmetic average of the single realization exponents, $\overline{\alpha}=N_{r}^{-1}\sum_{r=1}^{N_{r}}\alpha_{r}$. Even though the latter is not correct, the fact that both yield very similar values attests to the robustness of the average exponent, notwithstanding sizeable fluctuations. We estimate the scaling exponent $\alpha$ in the interval $[2.7-33.1]$ mm as in Deco2020 and we obtained $\alpha\approx 2/5$ for $\delta=5.55$ mm, and $\alpha\approx 2/3$ for $\delta=5.88$ mm.
IV.1 Scaling exponents as a function of the decay length
Figure 3: A) The blue line shows the average slope of $\alpha=\log S(d)/logd$ for different $\delta$s, given the connectivity matrix $J_{ij}=\exp{(-d_{i,j}/\delta)}$. The horizontal dotted lines correspond to $\alpha_{RW}=1$ and $\alpha_{T}=2/3$, corresponding to random walk and homogeneous incompressible turbulence, respectively, while the dot-dashed horizontal line corresponds to Deco et al. “brain” value $\alpha_{D}=1/2$. The orange line shows the function $\alpha(\delta)$ with a connectivity matrix obtained by randomly shuffling all the weights $J_{ij}$, which results into a basically uncorrelated signal, indicating the importance of the metric structure of the weights. The plot stops at $\delta=10$, a region where the signal is highly correlated ($\alpha\sim 1.4$), but still below the smooth regime marked by $\alpha=2$. A visual extrapolation indicates that the scaling regime would remain non-smooth ($\alpha<2$) even in the limit of an ”infinite-brain”, $\delta\to\infty$), the global brain ($d\sim 25\delta$) being pretty close to the infinite-brain limit. Note however that in this region $B(d)\sim B(0)$, so that the numerical calculation of the exponent becomes less and less accurate (a fully correlated signal with $B(d)=B(0)$ yields formally $\alpha\to\infty$.) Conversely, a brain below $4$ mm, would exhibit ”chaotic” behaviour, with scaling exponents close to zero, i.e. no scaling regime. The bottom panel shows the neuron activity mapped on the Schaefer’s cerebral cortical parcellation atlas coordinates for (B) $\delta=4$ (“chaotic”), (C) $\delta=5.99$, and (D) $\delta=6.66$ (random walk), respectively.
We repeated the previously depicted procedure for a sequence of values of the decay length in the range $[3-10]$ mm and measured the associated scaling exponents. The values of $\alpha(d)$ for different realizations are shown in Fig.3.

### v2

@@
Evidence of Scaling Regimes in the Hopfield Dynamics of Whole Brain Model
Giorgio Gosti,1,2∗ Sauro Succi,1,3 Giancarlo Ruocco1
1Center for Life Nano- and Neuro-Science, Istituto Italiano di Tecnologia,
Viale Regina Elena 201, I-00161, Rome, Italy
2Istituto di Scienze del Patrimonio Culturale, Consiglio Nazionale delle Ricerche,
00010, Montelibretti (RM), Italy
3Istituto per le Applicazioni del Calcolo del Consiglio Nazionale delle Ricerche,
via dei Taurini 19, I-00185, Rome, Italy
4Dipartimento di Fisica, Università di Roma “La Sapienza”,
P.le Aldo Moro 5, I-00185, Roma, Italy
To whom correspondence should be addressed; E-mail: giorgio.gosti@cnr.it
@@
It is shown that a Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently introduced by Deco et al., according to which the process of information transfer within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows, although with a more singular exponent, $1/2$ instead of $2/3$. Both models employ a coupling strength which decays exponentially with the Euclidean distance between the nodes, but their mathematical nature is very different, Hopf oscillators versus a Hopfield neural network, respectively. Hence, their convergence for the same data parameters, suggests an intriguing robustness of the scaling picture. The present analysis shows that the Hopfield model brain remains functional by removing links above about five decay lengths, corresponding to about one sixth of the size of the global brain. This suggests that, in terms of connectivity decay length, the Hopfield brain functions in a sort of intermediate “turbulent liquid”-like state, whose essential connections are the intermediate ones between the connectivity decay length and the global brain size. Finally, the scaling exponents are shown to be highly sensitive to the value of the decay length, as well as to the number of brain parcels employed. As a result, any quantitative assessment regarding the specific nature of the scaling regime must be taken with great caution.
Introduction
@@
In a recent paper  (?), it was argued that the information transfer within the human brain may proceed in close analogy with the mechanisms which govern mass and energy transport in turbulent fluids. The analogy refers to the emergence of correlated spacetime patterns which would facilitate and possibly optimize information transfer across the brain, in the “same” way as fluid turbulence enhances energy transport across fluids. This is interesting in many respects, not least the prospects for future neurocomputers  (?).
Based on an extended dynamical system of oscillators (Hopf model), supplemented by large-scale neuro-imaging empirical data on the connections between different brain areas, the authors of  (?) numerically showed that the statistical correlations of human brain signals may exhibit a scaling regime similar to the one observed in turbulent fluids.
@@
At a closer scrutiny, though, the analogy does not appear to be quantitative, meaning that the spatial scaling exponents of the second order structure function $S_{2}(d)$=$\langle(O(d)-O(0))^{2}\rangle$$\sim d^{\alpha}$, $O$ being a suitable order parameter, are not the same, about $\alpha$$\sim$1/2 for the Hopf model of the human brain against $\alpha$=2/3 for fluid turbulence. Hence “brain turbulence” appears to be more singular than fluid turbulence.
Yet, a transient scaling regime in space is found in either cases, which is intriguing per se, regardless of the specific value of $\alpha$.
@@
In this paper, we examine the robustness of the aforementioned analogy by conducting a similar analysis based on the same neuroimaging data for brain connectivity, but using a very different model: a Hopfield neural network. Additionally, we measure the scaling exponents as a function of the connectivity decay length ($\delta$), discovering a much richer scenario, including a dependence on the number of parcels (network nodes) ($N$), to be detailed later in the paper.
@@
The recurrent Hopfield neural network model is a well-established model for small-scale neuronal networks  (?, ?, ?). It is a neural network model based on binary McCulloch-Pitts neurons  (?, ?, ?, ?) in which every processing unit ($i$) is connected to all the other ones ($j$) through a set of weights ($J_{ij}$). Most of the literature considers fully connected and symmetric Hopfield networks since this simplifies the mathematical analysis of the model, although recently the investigation of diluted networks  (?, ?, ?, ?, ?) and asymmetric networks  (?, ?, ?, ?, ?, ?, ?, ?) was introduced.
@@
Usually, a recurrent Hopfield neural network is trained to store patterns or memories as steady-state attractors with the Hebbian prescription. In this case, the coupling matrix $J_{ij}$ takes the form of the sum of diadiics, and therefore it is symmetric and fully connected. The maximum storage capacity of such a network is found to scale linearly with the number of nodes $N$. The capacity can be increased by the process of learning (Hebbian learning) whereby the simultaneous activation of two nodes reinforces their connection. This is inspired by biological neural networks where the simultaneous activation of neurons leads to increments in synaptic strength  (?), and it allows the network to store a larger number of patterns  (?). Yet, this quantity ($C$) still remains linear with $N$: $C\sim N$, being the proportionally factor of the order of unity. Lately, researchers have also investigated how unlearning methods can improve the stability of the stored patterns  (?, ?, ?).
@@
Interestingly, it has been recently demonstrated both numerically  (?, ?) and theoretically  (?, ?) that random, non-Hebbian, coupling matrices have a storage capacity much larger than the Hebbian ones. Indeed the storage capacity $C$ of random matrices scales exponentially with $N$: $C\sim e^{\Sigma N}$ being $\Sigma$ the complexity. The latter quantity is found to be maximized for asymmetric and diluted matrices, with values that match the connectivity found in the hippocampus areas.
@@
The approach outlined in reference  (?) diverges from both the Hebbian and random connectivity models. It relies on a coupling matrix, which is derived from experimental observations of the connections between human brain regions. This approach assumes that the coupling strengths are all positive, deterministic, and specifically proportional to a given function of the distances between nodes. $J_{ij}=e^{-d_{ij}/\delta}$, $d_{ij}$ being the Euclidean distance between nodes $i$ and $j$ in the network, and $\delta$ a decay length, which, in  (?) are kept fix to match the empirical HCP dMRI tractography of the human brain: $\delta=5.55$ $\mathrm{m}\mathrm{m}$.
At variance with the latter approach, where the dynamics is based on the Hopf oscillator model, here we use the same coupling matrix of  (?), but employ the Hopfield dynamics.
The use of Hopfield networks to investigate the collective dynamics of the brain activity is not new. Indeed,  (?), and  (?) used Hopfield models (or Ising-Spin models with Glauber dynamics) to study the complex landscape of brain activation dynamics. These schemes do not interpret the nodes of the Hopfield network as single spiking neurons but rather as aggregates of neural mass (parcel), namely mesoscale brain regions switching between active and rest states. In the active states many neurons simultaneously fire while in the rest states most neurons are silent. These models do not aim at establishing a rigorous connection between the microscopic description of the brain and a corresponding mesoscopic one, but posit instead a physically-informed heuristic picture of the collective brain dynamics.
Broadly speaking, the Hopf model is based on the idea of reproducing the empirically observed information transfer between areas that takes place through (partial) synchronization of collective oscillations. In this framework, the natural collective variables describing the global brain dynamics are mesoscale oscillators obeying the Hopf dynamics.
The Hopfield model used in this work draws inspiration from a very different idea, namely to carry over the discontinuous divide between active and rest states from the microscale level of single neurons to the mesoscale level of neural parcels, as discussed above. Formally, this amounts to assuming that coarse-graining of the Hopfield model would leave the model invariant but with a different set of weights, in our case an exponential decay. Such model proves useful to show that the dynamical repertoire of brain activity composed of sudden avalanches and bursts, can be described in terms of critical systems perturbed by noise  (?, ?).
Due to the significant difference between the fundamental assumptions behind these two models, inspecting their commonolaties, as well as their points of departure, starting from the same empirical connectivity information, is of decided interest to improve the mesoscale description of the global brain dynamics,
@@
ii) At the $\delta$ value investigated in  (?) ($\sim 5.5$ mm) and for $N=1000$, we found $\alpha\sim 2/5$, close to the one reported by Deco et al.  (?) [^1] In Ref.  (?) the used parameter is the inverse of the decay scale, $\lambda=1/\delta$.;
iii) For $N=1000$ the scaling exponent $\alpha$ shows a sigmoid-like trend with the decay length $\delta$. Importantly, for $N=1000$, the functional relation $\alpha(\delta)$ shows a steep increase in the region around $\delta\sim 5.55$ mm, the value indicated in  (?) as the physiologically relevant one from the analysis of empirical HCP dMRI tractography with a $N=1000$ parcellation. This means that the specific value of the decay length has a significant impact on the scaling exponent; for instance, a slight increase to $\delta$=5.88 mm would yield a scaling exponent slightly above $2/3$, the value associated with homogenoeus incompressible turbulence. Conversely, a slight decrease to $\delta$=5 mm, would yield $\alpha$$\sim$0.1, a much less organized regime.
iv) the sigmoid dependence of $\alpha$ on $\delta$ changes with $N$, becoming steeper and narrower at increasing $N$.
@@
Discrete-time Hopfield Recurrent Neural Networks
Following  (?, ?, ?, ?, ?, ?, ?, ?), we consider a network of $N$-binary nodes that can be either firing or silent.The state of each node is a binary variable $s_{i}$, with $i=1,2,\ldots,N$, such that $s_{i}=-1$ or $s_{i}=1$, so that the network state is represented by a binary string $\mathbf{s}=(s_{0},s_{1},\ldots,s_{N})$. Neurons interact through the connectivity matrix $\mathbf{J}$, with matrix elements $J_{ij}$ following the same exponential dependence $J_{ij}=e^{-d_{ij}/\delta}$ used in  (?). We assume a discrete time $t$, and, at each time step $t$, the evolution of the neuron state $s_{i}(t)$ is given by the following non-linear dynamic equation,
@@
where the activation function $sign(x)$ is defined to be 1 for $x$$\geq$0, and -1 otherwise. Consequently, whenever the summation of the inputs on node $i$ is above zero, the node is active (“fires”), otherwise, it remains silent.
We determined the couplings matrix $J_{ij}$ using Schaefer’s cerebral cortical parcellation atlas as in  (?). This parcellation was obtained from the analysis of a large dataset of neural activity dynamics with a sample size of 1489  (?). This dataset is publicly available, we used Schaefer’s parcellations with $N$ ranging from $200$ to $1000$. The Schaefer’s parcel centroids are mapped onto the MNI152 volumetric space [^2] Github repository dataset url  (?): https://github.com/ThomasYeoLab/CBIG/tree/master/stable\_projects/brain\_parcellation/Schaefer2018\_LocalGlobal/Parcellations/MNI/Centroid\_coordinates. The code for reproducing the simulations can be found in the git hub repository https://github.com/ggosti/HopBrain.
The Structure Factor $S_{2}(d)$
We used  (?) adaptation of Kolmogorov’s concept of structure functions. In turbulence, the Kolmogorov’s structure factor applies to the longitudinal velocity but in our case it represents the activity of the parcel as measured through the BOLD signal in a fMRI experiment.
To compute the structure factor we use the auxiliary covariance function $B(d)$ defined as:
$B(d)=\sum_{ij}\frac{1}{{\cal{N}}(d)}\sum_{d_{ij}\in d}s_{i}s_{j}$ (2)
@@
The function $B(d)$ relates to the structure factor as follows  (?):
$S_{2}(d)$ $=$ $2\left[B(0)-B(d)\right]$ (5)
@@
The Hopfield dynamics is deterministic, and since the phase space of the nodes’ states is finite, the asymptotic trajectories must necessarily be periodic. Therefore, as also discussed in  (?), Hopfield Recurrent Neural Networks, given an initial state, consistently end up into a specific steady state (“attractor”). The steady states are deterministic attractors composed of either limit cycles formed by repeating sequence of states, or a single stationary state. Given that in our case the couplings $J_{ij}$ are symmetric, we expected theoretically, and experimentally confirmed that the steady states are always composed of a single fixed point. In the absence of a further external input (“external stimuli”), and given sufficient time, the network gets stuck in an attractor and the whole activity is constant in time, hence no time summation was taken in the previous equations. This is of course a very crude representation of the brain dynamics, which receives continuously “internal” and “external” inputs, that perturb the system to constantly explore a landscape shaped by the collection of steady states. Indeed,  (?), and  (?) show how the whole set of attractors allows us to reproduce the dynamical repertoire and the functional correlations of the brain activation dynamics.
@@
Given different values of the decay length $\delta$, we have run the Hopfied model for a system composed by $N=1000$ parcels (with pairwise distances consistent with the Schaefer’s cerebral cortical parcellation atlas  (?)). The structure functions have then been calculated by following the system until the fixed points were reached, and by averaging over $N_{r}$=1000 randomly chosen initial states. As examples, figure 1 shows the structure factor $S_{2}(d)$ for respectively $\delta$=5.55 mm (Fig. 1A) and $\delta$=5.88 mm (Fig. 2B). The black points with the error bars represent the average $S(d)$ computed over 1000 different realizations. The dashed green line corresponds to $\alpha$=2/3, the orange line corresponds to $\alpha$=1/2, and the blue line corresponds to $\alpha$=2/5. The $\alpha$=1/2 value (orange line) is the value estimated in  (?). The $\alpha$=2/3 is the exponent associated with a turbulence. Due to the significant level of statistical fluctuations, we have estimated the average exponent in two different ways, first as the correct ensemble average $\langle S_{2}(d)\rangle$$\sim d^{\alpha}$ and also as an arithmetic average of the single realization exponents, $\overline{\alpha}$=$N_{r}^{-1}\sum_{r=1}^{N_{r}}\alpha_{r}$. Even though the latter is not correct, the fact that both yield very similar values attests to the robustness of the average exponent, notwithstanding sizeable fluctuations. For consistency with  (?), we estimate the scaling exponent $\alpha$ fitting the points in figure 1 in the interval 2.7 - 33.1 mm and we obtained $\alpha\approx$2/5 for $\delta$=5.55 mm, and $\alpha\approx$2/3 for $\delta$=5.88 mm.
Scaling exponents vs. the decay length
We repeated the previously depicted procedure for a sequence of values of the decay length in the range $[3-10]$ mm and measured the associated scaling exponents. The values of $\alpha(d)$ for different realizations are shown in Fig. 2 as blue points. The code stand instructions to reproduce the simulation and the 3D model files (gltf) are available at the repository https://github.com/ggosti/HopBrain. This code also shows how to generate 3D maps for the Schaefer parcellation’s nodes activation states inspection, and how to export gltf files that can be viewed in immersive environments using open source webXR apps such as Aton which are part of the H2IOSC open cloud  (?, ?) (Fig. 2B-D).